OpenStudy (study_buddy99):

evaluate 64^-2/3 (Alg II. I award medals)

OpenStudy (astrophysics):

Hint: \[\huge x^{-n} = \frac{ 1 }{ x^n }\]

OpenStudy (study_buddy99):

\[\frac{ 1 }{ 64^{2/3} } ???\]

OpenStudy (anonymous):

Yeah, can you keep going?

OpenStudy (anonymous):

First of all, remember to factor 64.

OpenStudy (anonymous):

\[ 64=2\times 32\\ 32 = 2\times 16 \\ 16 = 2\times 8\\ \ldots \]

OpenStudy (study_buddy99):

okay so 2 & 32... do I factor further?

OpenStudy (fortytherapper):

You could also use the rule that for x^m/n is the same as the n (root of x)^m So for 64^2/3, it would be (3(sqrt 64))^2 So you would find the cube root of 64, then square it

OpenStudy (study_buddy99):

... but it's negative

OpenStudy (fortytherapper):

Not when you use Astrophysic's formula

OpenStudy (study_buddy99):

but the answer is 1/16 and that would be 1/27

OpenStudy (priyar):

\[\frac{ 1 }{ 64^{2/3} } =\frac{ 1 }{(\sqrt[3]{64})^2 } = \frac{ 1 }{ (4)^2 } =?\]

OpenStudy (priyar):

any doubts?

OpenStudy (priyar):

@study_buddy99 hope u understood.. so what will be the answer?

OpenStudy (study_buddy99):

16

OpenStudy (priyar):

u mean" 1/ 16 "! thats correct!

OpenStudy (study_buddy99):

yay! Thank you!